Evaluating cost-cutting proposal

We are considering automating some part of an existing production process. The necessary equipment cost $80,000 to buy and install. That automation will save $22,000 per year (before taxes) by reduing labor and material cost. For simplicity, assuming that the equipment has a five-year life and is depreciated to zero on a straight-line basis over that period. It will actually be worth $20,000 in fives years. Should we automate? The tax rate is 34 %, and the discount rate is 10 %.
Answer #1
1] INITIAL INVESTMENT:
Cost of the new machine $         80,000
Initial investment $         80,000
2} OPERATING CASH FLOWS:
Annual savings in material & labor costs $         22,000
-Depreciation = 80000/5 = $         16,000
=NOI $ 6,000
-Tax at 34% $ 2,040
=NOPAT $ 3,960
+Depreciation $         16,000
=Annual OCF $         19,960
3] After tax salvage value = 20000*(1-34%) = $         13,200
Terminal non operating cash flow $         13,200
4] PV of annual OCF = 19960*(1.1^5-1)/(0.1*1.1^5) = $         75,664
PV terminal cash flow = 13200/1.1^5 = $ 8,196
PV of cash inflows $         83,860
Less: Initial investment $         80,000
NPV $ 3,860
5] As the NPV of the automation is positive, we
should automate.